OpenStudy (anonymous):

Correct me :D Which of the following is the slope between the points (-6, 0) and (2, -9)? M=y^2-y^1/x^2-x^1 M= -9-0/2-6 = -9/8 ?

OpenStudy (shamil98):

-9 - 0 / 2-(-6) = -9/8 You did it correctly.

OpenStudy (shamil98):

but..

OpenStudy (shamil98):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 }\] not y^2 - y^1/x^2-x^1

OpenStudy (anonymous):

Thank you :D